Subject: APO 3.5m users committee, minutes for October meeting

From: Suzanne Hawley

Submitted: Thu, 13 Oct 2011 15:41:12 -0700 (PDT)

Message number: 1250 (previous: 1249, next: 1251 up: Index)

NOTE: Science highlights for the ARC Board meeting are due by Monday, Oct 17, 
2011. Please forward to your users committee representative.

 		APO 3.5-m Users Committee Phonecon, 10/03/11

Attending:  Suzanne Hawley, Mark Klaene, Bill Ketzeback, Remy Indebetouw, Jon 
Holtzman, John Bally, Ed Turner (for Michael Strauss), Scott Anderson, Sean 
Moran, and Bruce Gillespie

Absent:  Al Harper


User feedback and comments from institutional representatives:

o Washington (Scott Anderson) - Scott had nothing to report.

o Colorado (John Bally) - John had nothing to report.

o Johns Hopkins (Sean Moran) - Sean had nothing to report.

o New Mexico State (Jon Holtzman) - Jon had nothing to report.

o Princeton (Ed Turner) - Ed had nothing to report.

o Virginia (Remy Indebetouw) - Remy had nothing to report.

o Chicago (Al Harper) - No report.


Discussion of telescope/instruments report:

Mark highlighted the period by stating the telescope and most instruments are 
working well.  Alaina Sheldon has replaced Joe Huenerhoff as a 3.5-m Observing 
Specialist.  The replacement B400 grating for DIS has been ordered, and a new 
slit-view camera for the echelle is also on order. For NIC-FPS, the recent 
problem with left-right balance seems to be fixed by replacing a video card. 
The NIC-FPS ICC also checked out OK, and the residual read-out problems may be 
the result of ice buildup on a new connector that was installed during the 
summer shutdown.  John Bally is now in charge of NIC-FPS engineering at CU.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                 9/8/11 through 9/27/11

0) Overview

We have an instrument PDR scheduled for mid October, an NMSU class trip 
scheduled at the end of October and a UVa class trip scheduled for mid November 
so housing will be tight during those periods.  The Alamo/El Paso Shuttle has 
changed owners.  Rates and schedule have changed but their contact information 
remains the same.  Monsoonal moisture has been with us much of the month but we 
are starting to see a drying trend.  We are only about 70% of our normal 
rainfall for the year and expect a dry winter.

1) Telescope

The telescope has been fully operational with no problems reported.  The damage 
to the fire alarm system has been repaired. Obs Spec Joe Huehnerhoff is 
transitioning to day engineering crew to work on the new imaging camera.  He is 
being replaced on the observing crew with Alaina Shelden who has been a SDSS 
observer for the past 2 years.  Alaina came to use from Univ. of Florida after 
working on the MARVELS instrument.

2) Instruments

DIS replacement Blue 400 grating has been ordered.  Slitviewer and spectrograph 
focus has been checked and are scheduled for adjustment this week.

TSpec, Agile and SPIcam operated with no new issues.  Echelle is operational. 
A decision on a new grating with higher throughput is expected this month.  A 
new slitviewer camera (Apogee Alta) has been ordered.  The new camera will have 
lower dark current (0.08 vs 0.8 e/pix/sec) and better QE (95% vs 60%).  Issues 
regarding wavelength calibration are being investigated.

NIC-FPS has been the issue of the month.  Staff are troubleshooting the 
instrument while learning the intricacies of the detector electronics. The 
left/right side subtraction issue was isolated to one of the Leach controller 
video cards.  The image corruption issue is still with us. We continue to find 
ice buildup on one of the detector bulkhead connectors. It appears the new 
cables installed at CU this summer are heavier gauge wiring and will need to be 
replaced.  The instrument is currently not available for science use.

GFP software development has made progress this period.  Dewar cleanout was 


Science highlights for Board meeting:

As we do each year, Suzanne asks the Users Committee members to provide 
institutional science highlights for inclusion on the annual report to the ARC 
Board of Governors.  Please provide Mark and Suzanne a 1-page powerpoint slide, 
or a figure plus text, by 17 October.

===> ACTION:  Users Committee members to provide Mark and Suzanne with 
institutional science highlights for the year, by 17 October, 1 slide or 
figures plus text.


ReSTAR update:

Suzanne said that we have been notified by NOAO that they are not planning to 
submit a proposal to the NSF for ReSTAR funding this fall.  This is due to the 
current funding climate at the NSF, but it is possible if things change a 
proposal could be submitted next spring.


Feedback on Target-of-Opportunity observations:

Suzanne said that we have been handling sporadic ToO requests over the past few 
years on an ad hoc basis, which has worked satisfactorily.  Now, an astronomer 
at Princeton has submitted a larger ToO program for the 3.5-m that could be 
invoked frequently.  Based on discussions and feedback from the users 
committee, and operational considerations, Suzanne has proposed a policy that 
will restrict ToOs (those that preempt scheduled observations) to one per 
month.  The time given to the ToO will be taken from that institution's 
allotment or from Director's Discretionary Time. The ToO notification must 
occur not later than 2 pm on the day of the observations, and the payback to 
the preempted observer will be made as soon as possible.  Note that there is a 
"protection" block on the observing proposal form that can be employed to block 
a ToO preemption. Users are encouraged to make use of this block if their 
science would be severely impacted by preemption.

Remy thought that the proposed policy is reasonable, and asked what the payback 
time would be for a ToO preemption -- he suggested that a half night would be 
appropriate.  Suzanne said that she intends to be flexible, and would grant a 
payback that suited the individual situation.  Jon asked if the one-per-month 
frequency was adequate for the science requirements. Ed said that even though 
the Princeton PI could use as much ToO time as possible, this level of access 
would be useful.  Ed added that any ToO policy will work some of the time, but 
there will be exception cases where the system breaks down.  He suggested that 
we could scatter Princeton time throughout the quarter and handle ToO incidents 
locally with internal time swaps.  Scott felt that the proposed policy seems 
fine, and Sean added that it would have little impact on JHU programs.  John 
also agreed that the policy sounded good.  Suzanne will write up a formalized 
version of the ToO policy.  This quarter, she will allow three ToOs, in 2-hour 
blocks, and payback will be with already scheduled Princeton 2 hour blocks, or 


New imager - Science Requirements discussion:

Suzanne announced that the PDR for the new imager will be held at APO on 20-21 
October.  Chris Stubbs wrote a white paper on the instrument that was 
distributed and discussed last month, and we are assembling science 
requirements that will be forwarded to the instrument team at Harvard. Suzanne 
then polled the committee for comments.

Sean said that the JHU community is general positive toward the new imager.  He 
asked about the FOV of the camera, and Bill said that it is a circular field 
with 12 arcminute diameter.  Hal Weaver wondered if a simple detector upgrade 
for SPIcam would be a cheaper, effective alternative.  Suzanne said that 
Cynthia Froning looked at this option as part of the imaging camera white paper 
process 1-2 years ago, and determined that it would not be cheap nor easy. 
Scott said that of the UW users polled, 5 were in favor and 3 were interested 
in preserving narrow-band imaging capability, but if SPIcam or GFP can be used 
for narrow-band imaging, that would be acceptable.  He added that there are 
users interested in planetary transits, with short exposures.  Bill said that 
the camera can provide short exposures, and of course Agile is already being 
used for fast photometry.  Jon asked about guiding, i.e., how does the existing 
off-axis guider play with this instrument, and is it possible that the new 
camera can do on-chip guiding with sufficiently short exposures.  Suzanne said 
that the guiding issues will be discussed at length at the PDR.  Remy asked if 
we remove the off-axis guider, how does this affect the other instruments. 
Mark said that we are looking at a mechanical modification to the off-axis 
guider that would allow it to be withdrawn or inserted into the focal plane, 
depending on which instrument is mounted.  He pointed out that only the imagers 
on the telescope use the guider; the spectrographs guide off their slits.  In 
general, people felt that the users will like the new imager, but narrow-band 
capability with some instrument at APO is important.

Jon said that in a sense, the new imager is pretty much "defined".  One 
disadvantage is the filter set is fixed, and precludes people doing UBVRI 
imaging vs. Sloan colors.  Also, we won't have u-band imaging. Furthermore, 
scattered light and photometric accuracy are unknowns and will be important to 
many users.  Lastly, is the internal atmospheric dispersion correction really 
necessary?  Generally, we need to be thinking about future users and science 

Suzanne closed by saying that she feels that this imager is generally 
attractive to our present and future user community, and may be a unique 
facility for mid-sized telescopes worldwide.  Stray light and guiding are 
important issues to investigate, and she is still interested in any further 
community feedback prior to the PDR -- send comments to her and Mark.



[open from previous meeting]:

===> ACTION:  Committee members should forward the Stubbs white paper to their 
users, and solicit input for the camera's Science Requirements Document, 
including ranges of science programs desired.  Due by next Users Committee 
meeting (3 October).

      STATUS:  Discussed at this meeting, closed.

===> ACTION:  Committee members should discuss options and desiderata with 
their users for a new policy for handling ToO requests on the telescope. The 
summaries of these discussions should be forwarded to Suzanne.

      STATUS:  Discussed at this meeting, closed.

[new actions from this meeting]:

===> ACTION:  Users Committee members to provide Mark and Suzanne with 
institutional science highlights for the year, by 17 October, 1 slide or 
figures plus text for each highlight.


Next meeting:  The next Users Committee phonecon will be on Monday 7 November 
at 8:30 AM Pacific Time.  The agenda and other materials will be sent to the 
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1250 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to